Paul's age is six less than twice Lucien's age. Two years ago, the sum of their age was 29. How old is Lucien?
Answer by feliz1965(151)

Paul's age now = 2x - 6
Lucien's age now = x
Paul's age 2 years ago = (2x - 6) - 2
Lucien's age 2 years ago = x - 2
(2x - 6) - 2 + x - 2 = 29
3x - 10 = 29
3x = 10 + 29
3x = 39
x = 39/3
x = 13
Lucien is 13 years old.